State the horizontal asymptote of the rational function.


f(x)=x^2+6x-8/x-8
1
Expert's answer
2014-06-11T03:10:06-0400
Given two polynomial: h(x)=x^2+6x-8 (second degree) and g(x)=x-8 (first degree). As degree of f(x) is more than degree of g(x), we can determine that there is no horizontal asymptote.